# Knowledge Distillation with Structured Chain-of-Thought for Text-to-SQL

### v1

@@ Abstract
Deploying accurate Text-to-SQL systems at the enterprise level faces a difficult trilemma involving cost, security and performance. Current solutions force enterprises to choose between expensive, proprietary Large Language Models (LLMs) and low-performing Small Language Models (SLMs). Efforts to improve SLMs often rely on distilling reasoning from large LLMs using unstructured Chain-of-Thought (CoT) traces, a process that remains inherently ambiguous. Instead, we hypothesize that a formal, structured reasoning representation provides a clearer, more reliable teaching signal, as the Text-to-SQL task requires explicit and precise logical steps. To evaluate this hypothesis, we propose Struct-SQL, a novel Knowledge Distillation (KD) framework that trains an SLM to emulate a powerful large LLM. Consequently, we adopt a query execution plan as a formal blueprint to derive this structured reasoning. Our SLM, distilled with structured CoT, achieves an absolute improvement of 8.1% over an unstructured CoT distillation baseline. A detailed error analysis reveals that a key factor in this gain is a marked reduction in syntactic errors. This demonstrates that teaching a model to reason using a structured logical blueprint is beneficial for reliable SQL generation in SLMs.
@@ 1 Introduction
Text-to-SQL (NL2SQL or text2sql) has the potential to democratize data access [fei2014constructing]. The field has seen substantial performance advancements driven by the advent of Large Language Models (LLMs) [shi2025asurvey]. These models enhance the capabilities of natural language interfaces for databases by automatically translating natural-language user questions into SQL queries [hong2024next]. Nevertheless, widespread adoption in enterprises remains challenging due to a difficult trade-off among three interdependent factors: cost, security, and performance. This challenge can be understood as an Adoption Trilemma.
@@
Cost: High-performing models typically require significant computational resources, leading to high operational costs, whether using proprietary APIs or privately hosted LLMs [shi2025asurvey].
@@
Security: Relying on external APIs raises prominent security concerns, as transmitting potentially sensitive database schemas and sample records to third-party providers is often unacceptable in enterprise settings [hoang-etal-2025-distill].
@@
Performance: Selecting open-source models for local, private deployment to address cost and security issues often leads to the use of Small Language Models (SLMs), which typically lack adequate zero-shot accuracy for complex real-world queries [li2024can].
@@
To identify the limitations of current approaches, we examine the impact of recent advances in reasoning-driven prompting on Text-to-SQL performance. A notable portion of the recent performance gain in Text-to-SQL using LLMs can be attributed to in-context learning (ICL) [dong-etal-2024-survey]. ICL-based methods, particularly those that use decomposition and multi-step reasoning, have demonstrated substantial gains in performance on the Text-to-SQL task [hong2024next]. A prominent ICL technique for this is Chain-of-Thought (CoT), which encourages models to think step-by-step [wei2022chain, liu2025uncovering]. Such decomposition strategies are beneficial for generating complex SQL queries. For instance, DAIL-SQL [gao2024dail] enhances CoT through context-aware examples, while DIN-SQL [pourreza2023dinsql] and Divide-and-Conquer CoT (DC-CoT) [pourreza2024chase] improve accuracy by breaking questions into intermediate sub-queries. Building on this logical foundation, the Query Plan CoT (QP-CoT) guides the model through steps that mirror a database execution plan, simulating the logical process by which databases execute queries. By following this structured path, QP-CoT ensures that the SQL generation path aligns with the inherent logic of query execution [pourreza2024chase]. Although these reasoning techniques achieve considerable success, their effectiveness is observed almost exclusively with large LLMs. This reliance on large LLMs constitutes a significant limitation: these methods depend on the very models that exacerbate the cost and security challenges, thereby limiting their applicability in resource-constrained enterprise settings and contributing to the ‘Adoption Trilemma’.
@@
The performance limitations of SLMs, suitable for private deployment, are particularly pronounced in the Text-to-SQL domain. Standard benchmarks highlight a significant gap between large models and smaller open-source alternatives. For example, on the BIRD mini-dev benchmark, LLMs such as GPT-4o and Claude (150B parameters) achieve execution accuracies between 30% and 45%, whereas widely used SLMs such as Mistral, Mixtral and Qwen2.5 Coder (7B parameters) achieve only 4% to 12%111Performance on BIRD mini-dev, see https://github.com/bird-bench/mini_dev. This paper reveals that this severe performance degradation in SLMs persists even when advanced reasoning prompts are employed, despite their effectiveness in larger LLMs. This finding is consistent with recent literature, suggesting that large LLMs can adhere to the schema as long as the schema fits within their context window, a capability not observed in SLMs [maamari2024death]. In our experiments, the SLM (Qwen3-4B-Instruct-2507) prompted with QP-CoT fails primarily due to inadequate schema adherence, with a pronounced tendency toward schema hallucination, often generating non-existent tables or columns (see Figure 3(b)). Thus, while large LLMs benefit from structured reasoning strategies, SLMs do not internalize these logical decompositions. This breakdown in structured reasoning motivates an investigation into whether the structure of reasoning itself can be effectively transferred from large LLMs to SLMs.
@@
To bridge this performance gap, Knowledge Distillation (KD) [hinton2014distilling], which transfers reasoning ability from a capable teacher model, is a key strategy [shi2025asurvey]. KD is a model compression technique in which a smaller "student" model is trained to mimic the behavior of a larger, pretrained "teacher" model. Beyond compression, KD enables the transfer of complex reasoning [li2024explanations] and logical skills [hsieh2023cotdistilling, yao2023specializing]. Through KD, our aim is to build SLMs for Text-to-SQL that deliver accuracy comparable to larger LLMs while meeting enterprise cost and security requirements through private deployment. This work argues that KD is a promising approach in addressing this trilemma, but its effectiveness depends on the type of reasoning being distilled.
@@
Given that the core bottleneck for SLMs is replicating the teacher’s complex reasoning, the central question for KD is identifying the optimal form of the reasoning signal to distill. Building on the success of CoT prompting [wei2022chain], one approach is to distill the teacher’s natural language reasoning traces [li2024explanations]. This approach is exemplified by ReasonSQL, where the student model is trained on the teacher’s intermediate CoT steps in addition to the final SQL query. This unstructured approach provides a better knowledge signal than finetuning on SQL queries alone and has been shown to improve model accuracy [rossiello2025rationalization]. However, we argue that the structure of the reasoning signal itself is critical for effective distillation. We hypothesize that distilling knowledge using a more formal, structured representation of the reasoning process that directly reflects the logical steps of query execution could provide a clearer, less ambiguous, and more beneficial supervisory signal for distilling SLMs than unstructured CoT explanations.
@@
To evaluate this hypothesis, we introduce Struct-SQL, a framework for distilling structured reasoning. Figure 1 shows the KD workflow, contrasting the proposed structured reasoning approach against the standard unstructured CoT method. Within this framework, a state-of-the-art (SOTA) Teacher Model is used to generate a QP-CoT trace, which formally decomposes the query into a logical execution plan [pourreza2024chase]. This structured plan, together with the generated SQL query, constitutes the supervisory signal. A student model is then trained to replicate the entire structured output sequence (query plan, SQL). The formal query plan serves as a clear, hierarchical blueprint that guides the student model in learning the precise, logical steps of query construction, from schema linking and join-path selection to aggregation and filtering. During inference, this structured reasoning is retained: the student model is given the QP-CoT prompt to autonomously generate the query plan before synthesizing the final SQL query.
@@
We perform an extensive comparative analysis on the mini-dev BIRD benchmark [li2024can], evaluating our Struct-SQL framework against key baselines. Our contributions are as follows.
@@
We provide a comprehensive error analysis, showing that the improvement in Struct-SQL results from a marked reduction in syntactic errors (e.g., schema hallucination), demonstrating that the structured signal provides a clearer curriculum.
@@
We release the KD dataset of 1,300 structured reasoning traces, the Struct-SQL model, prompts and the code 222https://github.com/craterlabs/Struct-SQL-Distillation/.
@@ 2.2 Structured CoT via Query Execution Plan
Our structured CoT strategy is inspired by a database engine’s query execution plan. A query execution plan defines the precise sequence of steps that a database follows to access and manipulate data, often generated using an EXPLAIN command. We adopt this ICL prompt strategy from a recent work on a Query Execution Plan-based prompting strategy, QP-CoT [pourreza2024chase]. Instead of a free-form explanation, the Teacher Model is prompted to generate a query plan along with the final SQL query. As shown in Figure 2, this plan decomposes the query into a sequential execution flow that explicitly performs selections, filters, joins, and groupings through step-by-step table scanning and data manipulation, providing a systematic signal for the student model. Although the query plan has been explored as a prompting technique [pourreza2024chase], the key contribution here is to demonstrate the effectiveness of this structured reasoning as the primary teaching signal within a KD framework.
@@ 2.3 Experimental Setup
The experiments use the SQLite-based BIRD benchmark for the data [li2024can]. The BIRD training dataset is used for model training, while the BIRD mini-dev dataset serves for evaluation. The Teacher Model (), GPT-4o (OpenAI), is selected due to its leading performance on the BIRD mini-dev dataset and functions as an oracle to generate high-quality query plans. The Student Model () is Qwen3-4B-Instruct-2507 (Alibaba Cloud). It was chosen for its strong performance-to-size ratio, which enables low-latency, private deployment and thus addresses the Adoption Trilemma [yang2025qwen3]. All models utilize a single-pass inference mechanism and operate without multi-agent collaboration, self-consistency checks or external correction loops.
@@ 2.3.1 Model Configurations
Tuned Student Models: These configurations evaluate the Student Model after tuning on specific signals.
FN-Gold The Student Model was finetuned using the Gold SQL query () on the BIRD training dataset, with a basic system instruction that directly maps natural language to SQL [pourreza2023dinsql].
ReasonSQL (unstructured KD): The Student Model was finetuned on the complete Teacher sequence , where is the free-form CoT rationale [rossiello2025rationalization]. This configuration serves as the established baseline for evaluating the efficacy of standard unstructured reasoning distillation.
@@ 2.3.2 Distillation Dataset Construction
We then applied stratified, success-based sampling. The ID corpus is iterated using predefined SQL structure and syntax categories to achieve target distributions for the training set (1,000 samples) and the in-domain validation set (150 samples). These categories, as shown in Table 1, are differentiated by the complexity of the SQL structure, including single-table queries, queries with subqueries, queries with joins/set operations and queries that combine both joins/set operations and subqueries. For each candidate, inference was performed using GPT-4o. A sample is admitted only if the generated SQL is both syntactically valid and yields the correct result upon execution. New samples are generated until 1,000 successful training samples and 150 ID validation samples are obtained. The OOD validation set (150 samples) is constructed using the same sampling and validation process, but applied exclusively to the segregated OOD database pool to ensure that no schema overlaps. For both Struct-SQL and ReasonSQL, we used the same data generation pipeline to ensure a controlled, methodologically fair comparison. The datasets differ only in the format of the supervisory signal: unstructured CoT traces for ReasonSQL versus structured QP-CoT for Struct-SQL.
@@ 2.4.1 Post-Training Details
FN-Gold, ReasonSQL, and Struct-SQL were finetuned using Parameter-Efficient Fine-Tuning (PEFT) using Quantized Low-Rank Adaptation (QLoRA) [dettmers2023qlora] to improve training efficiency. PEFT was chosen for its proven effectiveness in adapting SLMs to new domains with small training datasets. Furthermore, PEFT has been shown to yield more stable models and mitigate the risk of catastrophic forgetting of foundational knowledge. QLoRA allows recovery of the full 16-bit finetuning task performance even when the base model is loaded at 4-bit precision [han2024parameterefficient]. The KD approach keeps the original model parameters fixed and only updates a small set of LoRA adapter parameters . We followed the QLoRA methodology [dettmers2023qlora] and applied adapters to all linear layers of the Transformer architecture. Based on preliminary testing on a subset of the development data, we selected and an alpha scaling of to ensure robust adaptation. Optimization was performed using AdamW with a learning rate of , a maximum input length of tokens, and a generation limit of tokens. We utilized a batch size of 6 for ReasonSQL and Struct-SQL, compared to 15 for the FN-Gold baseline on an NVIDIA H200 GPU. All models minimized completion loss. To balance in-domain accuracy with out-of-domain generalization, we used an early-stopping strategy (patience=8) that monitored the aggregated validation loss across both the ID and OOD validation sets.
@@ 2.4.2 Evaluation Metrics
The primary metric is Execution Accuracy (EX), which measures the percentage of generated SQL queries that execute without error and return the same result set as the ground-truth SQL query on the BIRD mini-dev set. A detailed analysis of model failure modes reveals why and where certain models perform best. To facilitate analysis, we categorize failures into three distinct types as defined in Table 2. This classification establishes a severity hierarchy for Text-to-SQL tasks: starting with the most severe Generation Failure (GEN), where the model produces no recognizable SQL output; followed by Syntactic Failure (SYN), indicating an unexecutable query due to grammar or schema errors; and finally, Logical Failure (SEM), representing a query that is syntactically correct but semantically inaccurate. Recognizing this progression is beneficial for discerning the specific challenges and improvements associated with each model.
@@
| Semantic Errors (SEM) | Incorrect Columns/Rows | Wrong columns/row count |
@@
| Empty Output | Empty result set | |
| Syntactic Errors (SYN) | No Such Column/Table | Non-existent table/column |
| Keyword Issue | Incorrect or misplaced SQL keyword | |
| Syntax/Clause Order | Incorrect clause order, missing parentheses, etc | |
| Generation Error (GEN) | N/A | No SQL output |

### v2

@@ Abstract
Deploying accurate Text-to-SQL systems at the enterprise level faces a difficult trilemma involving cost, security and performance. Current solutions force enterprises to choose between expensive, proprietary Large Language Models (LLMs) and low-performing Small Language Models (SLMs). Efforts to improve SLMs often rely on distilling reasoning from large LLMs using unstructured Chain-of-Thought (CoT) traces, a process that remains inherently ambiguous. Instead, we hypothesize that a formal, structured reasoning representation provides a clearer, more reliable teaching signal, as the Text-to-SQL task requires explicit and precise logical steps. To evaluate this hypothesis, we propose Struct-SQL, a novel Knowledge Distillation (KD) framework that trains an SLM to emulate a powerful large LLM. To implement this approach, we adopt the query execution plan as a formal blueprint to derive structured reasoning. Our SLM, distilled with a structured CoT, achieves an absolute improvement of 8.1% over an unstructured CoT distillation baseline. A detailed error analysis reveals that a key factor in this gain is a marked reduction in syntactic errors. This demonstrates that teaching a model to reason using a structured logical blueprint is beneficial for reliable SQL generation in SLMs.
@@ 1 Introduction
Text-to-SQL (NL2SQL or text2sql) has the potential to democratize data access [14]. The field has seen substantial performance advancements driven by the advent of Large Language Models (LLMs) [28]. These models enhance the capabilities of natural language interfaces for databases by automatically translating natural-language user questions into SQL queries [11]. Nevertheless, widespread adoption in enterprises remains challenging due to a difficult trade-off among three interdependent factors: cost, security, and performance. This challenge can be understood as an Adoption Trilemma.
@@
Cost: High-performing models typically require significant computational resources, leading to high operational costs, whether using proprietary APIs or privately hosted LLMs [28].
@@
Security: Relying on external APIs raises serious security concerns, as transmitting potentially sensitive database schemas and sample records to third-party providers is often unacceptable in enterprise settings [10].
@@
Performance: Selecting open-source models for local, private deployment to address cost and security issues often leads to the use of Small Language Models (SLMs), which typically lack adequate zero-shot accuracy for complex real-world queries [15].
@@
To identify the limitations of current approaches, we examine the impact of recent advances in reasoning-driven prompting on Text-to-SQL performance. A notable portion of the recent performance gains in Text-to-SQL using LLMs can be attributed to in-context learning (ICL) [3]. ICL-based methods, particularly those that use decomposition and multi-step reasoning, have demonstrated substantial gains in performance on the Text-to-SQL task [11]. A prominent ICL technique for this is Chain-of-Thought (CoT), which encourages models to think step-by-step [32, 19]. Such decomposition strategies are beneficial for generating complex SQL queries. For instance, DAIL-SQL [6] enhances CoT through context-aware examples, while DIN-SQL [23] and Divide-and-Conquer CoT (DC-CoT) [22] improve accuracy by breaking questions into intermediate sub-queries. Building on this logical foundation, the Query Plan CoT (QP-CoT) guides the model through the steps that mirror a database execution plan, simulating the logical process by which databases execute queries. By following this structured path, QP-CoT ensures that the SQL generation path aligns with the inherent logic of query execution [22]. Although these reasoning techniques achieve considerable success, their effectiveness is observed almost exclusively in the large LLMs. This reliance on large LLMs constitutes a significant limitation: these methods depend on the very models that exacerbate the cost and security challenges, thereby limiting their applicability to resource-constrained enterprise settings and contributing to the ‘Adoption Trilemma’.
@@
The performance limitations of SLMs, suitable for private deployment, are particularly pronounced in the Text-to-SQL domain. Standard benchmarks highlight a significant gap between large models and smaller open-source alternatives. For example, on the BIRD mini-dev benchmark, LLMs such as GPT-4o and Claude 3.7 Sonnet (150B parameters) achieve execution accuracies between 30% and 45% using standard prompting, whereas widely used SLMs such as Mistral, Mixtral and Qwen2.5 Coder (7B parameters) achieve only 4% to 12%111Performance on BIRD mini-dev, see https://github.com/bird-bench/mini_dev. This paper reveals that this severe performance degradation in SLMs persists even when advanced reasoning prompts are employed, despite their effectiveness in larger LLMs. This finding is consistent with recent literature, suggesting that large LLMs can adhere to the schema as long as the schema fits within their context window, a capability not observed in SLMs [21]. In our experiments, the SLM (Qwen3-4B-Instruct-2507) prompted with QP-CoT fails primarily due to inadequate schema adherence, with a pronounced tendency toward schema hallucination, often generating non-existent tables or columns (see Figure 3(b)). Thus, while large LLMs benefit from structured reasoning strategies, SLMs do not internalize these logical decompositions. This breakdown in structured reasoning motivates an investigation into whether the structure of reasoning itself can be effectively transferred from large LLMs to SLMs.
@@
To bridge this performance gap, Knowledge Distillation (KD) [9], which transfers reasoning ability from a capable teacher model, is a key strategy [28]. KD is a model compression technique in which a smaller "student" model is trained to mimic the behavior of a larger, pretrained "teacher" model. Beyond compression, KD enables the transfer of complex reasoning [16] and logical skills [12, 5]. Through KD, our aim is to build SLMs for Text-to-SQL that deliver accuracy comparable to larger LLMs while meeting enterprise cost and security requirements through private deployment. This work argues that KD is a promising approach to addressing this trilemma, but its effectiveness depends on the type of reasoning being distilled.
@@
Given that the core bottleneck for SLMs is replicating the teacher’s complex reasoning, the central question for KD is identifying the optimal form of the reasoning signal to distill. Building on the success of CoT prompting [32], one approach is to distill the teacher’s natural language reasoning traces [16]. This approach is exemplified by ReasonSQL, where the student model is trained on the teacher’s intermediate CoT steps in addition to the final SQL query. This unstructured approach provides a better knowledge signal than finetuning on SQL queries alone and has been shown to improve model accuracy [27]. However, we argue that the structure of the reasoning signal itself is critical for effective distillation. We hypothesize that distilling knowledge using a more formal, structured representation of the reasoning process that directly reflects the logical steps of query execution could provide a clearer, less ambiguous, and more beneficial supervisory signal for distilling SLMs than unstructured CoT explanations.
@@
To evaluate this hypothesis, we introduce Struct-SQL, a framework for distilling structured reasoning. Figure 1 shows the KD workflow, contrasting the proposed structured reasoning approach against the standard unstructured CoT method. Within this framework, a state-of-the-art (SOTA) Teacher Model is used to generate a QP-CoT trace, which formally decomposes the query into a logical execution plan [22]. This structured plan, together with the generated SQL query, constitutes the supervisory signal. A student model is then trained to replicate the entire structured output sequence (query plan, SQL). The formal query plan serves as a clear, hierarchical blueprint that guides the student model in learning the precise, logical steps of query construction, from schema linking and join-path selection to aggregation and filtering. During inference, this structured reasoning is retained: the student model is given the QP-CoT prompt to autonomously generate the query plan before synthesizing the final SQL query.
@@
To validate our hypothesis, we perform an extensive comparative analysis on the mini-dev BIRD benchmark [15], evaluating our Struct-SQL framework against key baselines. Our main contributions are as follows.
@@
We provide a comprehensive error analysis, showing that the improvement in Struct-SQL results from a marked reduction in the syntactic errors (e.g., schema hallucination), demonstrating that the structured signal provides a clearer curriculum.
@@
We release the code (github.com/craterlabs/Struct-SQL-Distillation) and, the model and the dataset (huggingface.co/collections/craterlabs/struct-sql) to facilitate reproducible research.
@@ 2.2 Structured CoT via Query Execution Plan
Our structured CoT strategy is inspired by a database engine’s query execution plan. A query execution plan defines the precise sequence of steps that a database follows to access and manipulate data, often generated using an EXPLAIN command. We adopt this ICL prompt strategy from a recent work on a Query Execution Plan-based prompting strategy, QP-CoT [22]. Instead of a free-form explanation, the Teacher Model is prompted to generate a query plan along with the final SQL query. As shown in Figure 2, this plan decomposes the query into a sequential execution flow that explicitly performs selections, filters, joins, and groupings through step-by-step table scanning and data manipulation, providing a systematic signal for the student model. Although the query plan has been explored as a prompting technique [22], the key contribution here is to demonstrate the effectiveness of this structured reasoning as the primary teaching signal within a KD framework.
@@ 2.3 Experimental Setup
The experiments use the SQLite-based BIRD benchmark for the data [15]. The BIRD training dataset is used for model training, while the BIRD mini-dev dataset serves for evaluation. The Teacher Model (), GPT-4o (OpenAI), is selected due to its leading performance on the BIRD mini-dev dataset and functions as an oracle to generate high-quality query plans. The Student Model () is Qwen3-4B-Instruct-2507 (Alibaba Cloud). It was chosen for its strong performance-to-size ratio, which enables low-latency, private deployment and thus addresses the Adoption Trilemma [33]. All models utilize a single-pass inference mechanism and operate without multi-agent collaboration, self-consistency checks or external correction loops. For a fair comparison with baselines, no external reasoning signals were provided at test time. During inference, the student model autonomously generates the Query Execution Plan based on the input prompt before generating the final SQL.
@@ 2.3.1 Model Configurations
Tuned Student Models: These configurations evaluate the Student Model after tuning it on specific signals.
FN-Gold The Student Model was finetuned using the Gold SQL query () on the BIRD training dataset, with a basic system instruction that directly maps natural language to SQL [23].
ReasonSQL (unstructured KD): The Student Model was finetuned on the complete Teacher sequence , where is the free-form CoT rationale [27]. This configuration serves as the established baseline for evaluating the efficacy of standard unstructured reasoning distillation.
@@ 2.3.2 Distillation Dataset Construction
We then applied stratified, success-based sampling, where samples are grouped by SQL complexity categories and only admitted if the Teacher Model generates syntactically valid and execution-correct SQL. The ID corpus is iterated using predefined SQL structure and syntax categories to achieve target distributions for the training set (1,000 samples) and the in-domain validation set (150 samples). These categories, as shown in Table 1, are differentiated by the complexity of the SQL structure, including single-table queries, queries with subqueries, queries with joins/set operations and queries that combine both joins/set operations and subqueries. For each candidate, inference was performed using GPT-4o. A sample is admitted only if the generated SQL is both syntactically valid and yields the correct result upon execution. New samples are generated until 1,000 successful training samples and 150 ID validation samples are obtained. The OOD validation set (150 samples) is constructed using the same sampling and validation process, but applied exclusively to the segregated OOD database pool to ensure that no schema overlaps. For both Struct-SQL and ReasonSQL, we used the same data generation pipeline to ensure a controlled, methodologically fair comparison. The datasets differ only in the format of the supervisory signal: unstructured CoT traces for ReasonSQL versus structured QP-CoT for Struct-SQL.
@@ 2.4.1 Post-Training Details
FN-Gold, ReasonSQL, and Struct-SQL were finetuned using Parameter-Efficient Fine-Tuning (PEFT) using Quantized Low-Rank Adaptation (QLoRA) [2] to improve training efficiency. PEFT was chosen for its proven effectiveness in adapting SLMs to new domains with small training datasets. Furthermore, PEFT has been shown to yield more stable models and mitigate the risk of catastrophic forgetting of foundational knowledge. QLoRA allows recovery of near full 16-bit finetuning task performance even when the base model is loaded at 4-bit precision [8]. The KD approach keeps the original model parameters fixed and only updates a small set of LoRA adapter parameters . We followed the QLoRA methodology [2] and applied adapters to all linear layers of the Transformer architecture. Based on preliminary testing on a subset of the development data, we selected and an alpha scaling of to ensure robust adaptation. Optimization was performed using AdamW with a learning rate of , a maximum input length of tokens, and a generation limit of tokens. We utilized a batch size of 6 for ReasonSQL and Struct-SQL, compared to 15 for the FN-Gold baseline on an NVIDIA H200 GPU. All models minimized completion loss. To balance in-domain accuracy with out-of-domain generalization, we used an early-stopping strategy (patience=8) that monitored the aggregated validation loss across both the ID and OOD validation sets.
@@ 2.4.2 Evaluation Metrics
The primary metric is Execution Accuracy (EX), which measures the percentage of generated SQL queries that execute without error and return the same result set as the ground-truth SQL query on the BIRD mini-dev set. A detailed analysis of model failure modes reveals why and where certain models perform best. To facilitate analysis, we categorize failures into three distinct types as defined in Table 2. This classification establishes a severity hierarchy for Text-to-SQL tasks: starting with the most severe Generation Failure (GEN), where the model produces no recognizable SQL output; followed by Syntactic Failure (SYN), which indicates an unexecutable query due to grammar or schema errors; and finally, Logical Failure (SEM), representing a query that is syntactically correct but semantically inaccurate. Recognizing this progression is beneficial for discerning the specific challenges and improvements associated with each model.
@@
| Generation (GEN) | — | No SQL output |
| Syntactic (SYN) | No Such Column | Non-existent column reference |
| No Such Table | Non-existent table reference | |
| Keyword Issue | Incorrect or misplaced SQL keyword | |
| Syntax/Clause Order | Incorrect clause order, missing parentheses | |
| Other | Unclassified syntactic errors | |
| Semantic (SEM) | Column Mismatch | Incorrect number of columns |
| Row Mismatch | Incorrect number of row | |
| Row & Column Mismatch | Incorrect number of columns and rows | |
@@
| Empty Output | Empty result set |